OpenStudy (anonymous):

8 1/8 divided by 2 1/4

OpenStudy (anonymous):

@slade

OpenStudy (anonymous):

3 11/18

OpenStudy (anonymous):

nope @slade

OpenStudy (anonymous):

u yes

OpenStudy (kash_thesmartguy):

First step: mix the fractions and find the reciprocal of the first fractio:\[\frac{ 65 }{ 8 }*\frac{ 9 }{ 4 }\]

OpenStudy (anonymous):

65/18 then

OpenStudy (anonymous):

@Kash_TheSmartGuy do ii multiply

OpenStudy (study_together):

First convert it into an improper fraction.\[\frac{ 65 }{ 8}\]\[\frac{ 9 }{ 4 }\]\[\frac{ 65 }{ 8 }\div \frac{ 9 }{ 4 }\]